OpenStudy (anonymous):

Ally is wearing a witch’s costume for Halloween. Her hat is in the shape of a cone with a diameter of 7 inches and a height of 18 inches. What is the volume of Ally’s hat? 924 in3 180 in3 231 in3 693 in3

OpenStudy (anonymous):

I had to look this one up: Volume of a cone as a function of radius and height is\[V(r,t)=\frac{pi r^2 h}{3}\]It would seem your radius is about 7/2 and height eighteen. You should be able to get a pretty good approximation.....

OpenStudy (anonymous):

so 7/2*18?

OpenStudy (anonymous):

Did you look at the formula?

OpenStudy (anonymous):

well what is pir o.o?

OpenStudy (anonymous):

pie*3.5*18/3?

OpenStudy (anonymous):

Having problem with the graphics program a little.... pi is about 22/7, the ratio between a circle's circumference and diameter. r is the radius, in the case of this problem, 7/2. We will need to square it. h is the height of the cone. In this problem, it is eighteen.

OpenStudy (anonymous):

ohhh pie and r...

OpenStudy (anonymous):

[22/7\times7/2^{2}\times18\div3=213 :D thank you\]

OpenStudy (anonymous):

I get closer to 230....

OpenStudy (anonymous):

sorry ment 231

OpenStudy (anonymous):

[22/7\times7/2^{2}\times18\div3=231 :D thank you\]